OpenStudy (anonymous):

what is the derivative of y = ((4-x)^2/x)? a. 3 -(16/x^2) b. 16x^2-3 c. 3x^2 - 16 d. (3/x^2)+6

OpenStudy (anonymous):

i edit it because its very confusing the answer is not here.

OpenStudy (anonymous):

y=(4x^2+2x)/x^2 => y = 4x^2 / x^2 + 2x / x^2 => y = 4 +...

OpenStudy (anonymous):

((4-x)(4-x))/x^2

OpenStudy (dumbcow):

you need to use quotient rule \[\frac{f}{g}' = \frac{f'g - fg'}{g^{2}}\] f = (4-x)^2 g = x f' use chain rule u = 4-x du = -1 d/du u^2 = 2u --> f' = -2(4-x) = -8 +2x g' = 1

OpenStudy (anonymous):

uh...i got a -16/x^2 +1 when i expand and did it....i dunno, jst said it coz it seemed sorta close to the answers

OpenStudy (anonymous):

yes mauve thats y im carefull all the answer are close that i wanted to choose letter E

OpenStudy (anonymous):

i dunno, i tried the product and the quotient both....bt ddnt come close enuf. anyways goodluck :-)

OpenStudy (dumbcow):

yes the correct answer is (x^2-16)/x^2 or 1 - 16/x^2

OpenStudy (anonymous):

ok im 100% sure i put letter E now